OpenStudy (anonymous):

the width of photograph is 4 centimeters more than three tenths of the length. if the width is 13cm. find the length?

OpenStudy (anonymous):

according to your question width is 13 cm so let length be x cm so w=4+3/10 x 13=4+3x/10 3x/10=9 3x will be 90 so x will be ? hope you got it@fifi4

OpenStudy (anonymous):

30 , ty !

OpenStudy (anonymous):

yuup you got it @fif4